# Consumption Partial Insurance

### v3

@@ A.2 Taylor approximation to lifetime budget constraint
$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})(\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1})^{2}=0$
@@ Appendix B Identification details
The moments of income measurement error are not separately identifiable from the moments of the transitory shock (Blundell et al., 2008). This is why we have assumed the income error to be Gaussian (but we let the consumption error unrestricted), so we specify
@@
Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks that comprise a two-year period. We thus assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t-1}}^{2}$, $\gamma_{\zeta_{t}}\approx\gamma_{\zeta_{t-1}}$, and $\kappa_{\zeta_{t}}\approx\kappa_{\zeta_{t-1}}$, which are exactly true if the underlying distribution is stationary. Similarly, we assume that $\sigma_{\xi_{t}}^{2}\approx\sigma_{\xi_{t-1}}^{2}$. For simplicity of the next statements, we also assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t+1}}^{2}\approx\sigma_{\zeta_{t+2}}^{2}$, $\sigma_{v_{t}}^{2}\approx\sigma_{v_{t-2}}^{2}\approx\sigma_{v_{t+2}}^{2}$, and $\sigma_{u^{y}_{t}}^{2}\approx\sigma_{u^{y}_{t-2}}^{2}\approx\sigma_{u^{y}_{t+2}}^{2}$. These last restrictions are not needed for identification with biennial data, but they simplify the illustration.
@@ B.1 Income process parameters
The second and third moments of shocks are then given by
@@ B.3 Quadratic consumption function parameters
$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})\\ \end{array}\right),$
@@
where the first inequality follows from the lower bound on the kurtosis. The term inside the squared brackets in the determinant is thus strictly positive, and $\text{det}(\mathbf{A})<0$. The matrix of coefficients is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli. [^2] If the distributions of shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$, so the determinant would not always be strictly away from zero.
@@
An interesting special case arises when the distributions of permanent and transitory shocks are normal and have the same variance, and there is no measurement error. In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}<0$ (where $\sigma_{t}$ is the dispersion of the generic distribution), $\phi_{t}^{(1)}=(1/(2\sigma^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$. Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of a typical permanent shock away from the mean, is separately identified from $\phi_{t}^{(1)}$. Therefore, identification does not require the shocks to be non-Gaussian. Despite normality in this case, tail shocks are still associated with a different pass-through from that of average shocks. Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
@@
Upon identification of the transmission parameters, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2})$ identifies $\sigma_{u^{c}_{t}}^{2}$, while ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ identifies $\sigma_{\xi_{t}}^{2}$. Higher moments of the consumption error or of $\xi$ involve higher than fourth-order moments of income, which we do not model.
@@ B.4 Targeted moments
There is a large number of over-identifying moments that we target in the estimation of the various specifications of the model. Table B.1 lists the moments targeted in the various specifications of the income process, while table B.2 lists the moments targeted in the different specifications of the consumption model. For simplicity, we assume stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters. These are not identifying assumptions but they make the illustration below more compact by allowing us to pool common terms together.
@@ B.4.2 Moments of consumption – linear function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+8(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+4(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}$
$+4(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+8(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+\kappa_{u^{c}}+3(\sigma_{u^{c}}^{2})^{2}$
@@ Appendix C Empirical details and additional results
C.1 Comparison with administrative moments
Table C.1 reports the empirical second, third, and fourth moments of the cross-sectional distribution of unexplained growth in male earnings (panel A) and household earnings (panel B), as opposed to the moments of household disposable income in table 2 in the paper. Column 1 presents moments from our baseline sample in the PSID, while column 2 presents the corresponding statistics, wherever available, from the U.S. Social Security Administration data reported in Guvenen et al. (2021).
@@ C.1 Comparison with administrative moments
Notes: The table presents the second, third, and fourth moments of earnings growth. Skewness and kurtosis correspond to the third and fourth standardized moments respectively. Column 1 reports moments of biennial earnings growth in our baseline sample in the PSID (we maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the PSID, this corresponds to a difference over two calendar years in this case), while column 2 reports moments of male earnings growth from the U.S. Social Security Administration data. The latter moments concern the one-year male earnings growth, unconditional on past earnings, and correspond to averages over ages 30-55 of, respectively, dispersion squared, skewness, and kurtosis, as reported in online appendix tables C23-C25 in Guvenen et al. (2021). Block bootstrap standard errors are in parentheses.
C.2 Sources of tail income shocks
Table C.2 provides evidence on the sources of large income reductions that contribute to income left skewness and excess kurtosis. Columns 1 & 2 report the empirical probability that unemployment, disability, bad health, etc, occur in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively. Column 3 reports the $p$-value of the difference between the two; in most cases, the bottom 10% (effectively households facing large income reductions) also experience unemployment, worsening health, large drops in hours, and precarious housing at much higher rates than the rest.

### v4

@@ A.2 Taylor approximation to lifetime budget constraint
$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})(\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1})^{2}=0.$
@@ Appendix B Identification details
The moments of income measurement error are not separately identifiable from the moments of the transitory shock. This is why we have assumed the income error to be Gaussian (but we let the consumption error unrestricted), so we specify
@@
Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks within a two-year period. We thus assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t-1}}^{2}$, $\gamma_{\zeta_{t}}\approx\gamma_{\zeta_{t-1}}$, and $\kappa_{\zeta_{t}}\approx\kappa_{\zeta_{t-1}}$, which are exactly true if the underlying distributions are stationary. Similarly, we assume that $\sigma_{\xi_{t}}^{2}\approx\sigma_{\xi_{t-1}}^{2}$. For simplicity of the next statements, we also assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t+1}}^{2}\approx\sigma_{\zeta_{t+2}}^{2}$, $\sigma_{v_{t}}^{2}\approx\sigma_{v_{t-2}}^{2}\approx\sigma_{v_{t+2}}^{2}$, and $\sigma_{u^{y}_{t}}^{2}\approx\sigma_{u^{y}_{t-2}}^{2}\approx\sigma_{u^{y}_{t+2}}^{2}$. These last restrictions are not needed for identification with biennial data, but they simplify the illustration.
@@ B.1 Income process parameters
The second and third moments of shocks are given by
@@ B.3 Quadratic consumption function parameters
$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})\\ \end{array}\right),$ (B.1)
@@
where the first inequality follows from the lower bound on the kurtosis. The term inside the squared brackets in the determinant is thus strictly positive, so $\text{det}(\mathbf{A})<0$. The matrix of coefficients is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli. [^2] If the distributions of both permanent and transitory shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$ and the determinant would not always be strictly away from zero.
How do different levels of skewness and kurtosis connect with the transmission parameters? Consider the transmission of permanent shocks, abstracting from transitory shocks (i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$). $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$, which corresponds to the first equation in (B.1). Absent any transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks. Suppose $\phi^{(1)}_{t}$ is fixed. We highlight three cases:
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}=0$. Letting $\gamma_{\zeta_{t}}<0$ as in the data, this equation restricts $\phi^{(2)}_{t}=0$. Tail shocks thus transmit similarly to non-tail ones, i.e., this configuration identifies a linear consumption process.
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$. Consumption growth and permanent shocks co-vary more strongly than in the linear consumption process, i.e., there is excess co-movement. For $\gamma_{\zeta_{t}}<0$, this configuration identifies $\phi^{(2)}_{t}<0$ and implies that negative shocks transmit more strongly than other shocks. Negative shocks are often the largest shocks (due to left skewness), which explains the excess co-movement. Ceteris paribus, a less skewed distribution of shocks ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement. If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$. Consumption growth and permanent shocks co-move less strongly than in the linear process, i.e., there is excess smoothness. For $\gamma_{\zeta_{t}}<0$, this identifies $\phi^{(2)}_{t}>0$ and suggests a muted pass-through of negative shocks. These are the largest shocks to income, so muting their pass-through suppresses the covariance between consumption and income growth. The opposite is true if $\gamma_{\zeta_{t}}>0$.
Another equation that pins down $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ is $2\gamma_{\zeta_{t}}\phi^{(1)}_{t}+2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}))\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$, which, abstracting from income measurement error, corresponds to the second line in (B.1). Absent transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and squared shocks so it captures the pass-through of shocks of large magnitude. We highlight again three cases (for $\phi^{(1)}_{t}$ fixed):
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}=0$. This identifies $\phi^{(2)}_{t}=0$, so tail shocks transmit similarly to non-tail ones. In fact, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})=2\gamma_{\zeta_{t}}\phi^{(1)}_{t}$ is directly implied by $\Delta^{2}c_{it}=\phi_{t}^{(1)}(\zeta_{it}+\zeta_{it-1})$, so this configuration identifies the linear consumption process.
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}<0$, which, for $\gamma_{\zeta_{t}}<0$, is true when ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$ is negative and large. This suggests that large shocks (of which negative shocks are the most likely) transmit strongly with a negative coefficient. It thus identifies $\phi^{(2)}_{t}<0$. Ceteris paribus, a less leptokurtic income distribution (lower $\kappa_{\zeta_{t}}$) turns $\phi^{(2)}_{t}$ more negative; a bell-shaped distribution is associated with fewer large shocks, thus necessitating a stronger transmission of said shocks to maintain a given excess co-movement.
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}>0$. This identifies $\phi^{(2)}_{t}>0$ for analogous reasons. For given excess co-movement, the smaller the kurtosis of permanent shocks, the larger the value of $\phi^{(2)}_{t}$ must be to achieve said excess co-movement.
In practice, $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are jointly pinned down by this and other equations of system (B.1), and they are also determined jointly with the transmission parameters of transitory shocks, to which analogous statements apply – skipped here for brevity. The sensitivity analysis is thus more nuanced so simulations can prove useful. Starting from the baseline estimates of the income process (table 2), reducing the left skewness of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by almost 10% (more negative). Independently reducing the kurtosis of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by 60% (more negative). The response of $\phi^{(1)}_{t}$ is much smaller, confirming that the values of skewness and kurtosis are crucial for identifying the pass-through of tail shocks.
@@
An interesting special case arises when the distributions of permanent and transitory shocks are normal, have the same variance between them, and there is no measurement error. In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the generic distribution. Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$. Therefore, identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks. Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks. Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
@@
Upon identification of the transmission parameters, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2})$ identifies $\sigma_{u^{c}_{t}}^{2}$, while ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ identifies $\sigma_{\xi_{t}}^{2}$. Higher moments of the consumption error or of the taste shock involve higher than fourth-order moments of income, which we do not model.
@@ B.4 Targeted moments
There is a large number of over-identifying moments that we target in the estimation of the various specifications of the model. Table B.1 lists the income moments targeted in each case, while table B.2 lists the moments targeted in the different specifications of the consumption model. For simplicity, we illustrate the next statements assuming stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters. These are not identifying assumptions but they make the illustration more compact by allowing us to pool common terms together.
@@ B.4.2 Moments of consumption – linear function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+8(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+4(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}+3(\sigma_{u^{c}}^{2})^{2}$
$+4(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+8(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}+\kappa_{u^{c}}$
@@ Appendix C Empirical details and additional results
C.1 Summary statistics in the PSID sample
Table C.1 presents summary statistics in the PSID sample over the period 1999–2019 (columns 1–2), as well as in subsamples sorted by wealth and education (columns 3–6).
Table C.1: Descriptive statistics, PSID
Low wealth High wealth
Baseline sample No college Some college No college Some college
mean median mean mean mean mean
(1) (2) (3) (4) (5) (6)
Consumption
Nondurable+services 26,766 22,895 21,280 24,566 23,675 31,479
Food at home 8,256 7,609 7,438 8,210 7,505 8,870
Utilities 3,515 3,238 3,351 3,205 3,471 3,787
Food out 3,318 2,511 2,387 2,928 3,109 4,037
Health 1,614 863 1,321 1,445 1,623 1,850
Transportation 5,866 4,900 5,283 5,489 5,711 6,406
Education 3,202 0 957 1,986 1,810 5,352
Childcare 977 0 546 1,259 537 1,136
Rent (or rent eq.) 18,105 13,821 8,910 12,631 16,449 26,020
Home insurance 894 747 585 712 908 1,142
Income
Disposable income 142,910 113,581 84,414 110,176 120,931 195,158
Taxable+transfer inc. 137,277 109,003 77,707 106,276 113,789 189,446
Earnings 118,423 97,154 71,504 99,798 90,950 158,616
Male 81,857 62,185 46,025 65,725 57,519 114,545
Female 36,567 29,183 25,479 34,074 33,430 44,070
Wealth 547,761 191,193 68,561 70,665 670,948 1,024,577
Demographics
% working (male) 0.89 1.00 0.89 0.94 0.82 0.89
% working (female) 0.79 1.00 0.79 0.82 0.78 0.78
Age (male) 46 45 44 42 50 48
% some college (male) 0.68 1.00 0.00 1.00 0.00 1.00
% some college (female) 0.71 1.00 0.42 0.79 0.48 0.87
Periods per household 5.63 5.00 4.90 4.88 5.86 6.62
Observations 20,866 4,258 5,159 2,428 9,021
C.2 Comparison with administrative moments
Table C.2 reports the empirical second, third, and fourth moments of the cross-sectional distribution of unexplained growth in male earnings (panel A) and household earnings (panel B), as opposed to the moments of household disposable income in table 1. Column 1 presents moments in the PSID sample, while column 2 presents the corresponding statistics, wherever available, from the U.S. Social Security Administration data of Guvenen et al. (2021).
@@ C.2 Comparison with administrative moments
Notes: The table presents the second, third, and fourth moments of earnings growth. Skewness and kurtosis correspond to the third and fourth standardized moments respectively. Column 1 reports moments of biennial earnings growth in the PSID (we maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the PSID, this corresponds to a difference over two calendar years), while column 2 reports moments of male earnings growth from the U.S. Social Security Administration data. The latter concern the one-year male earnings growth, unconditional on past earnings, and correspond to averages over ages 30-55 of, respectively, dispersion squared, skewness, and kurtosis, as reported in appendix tables C23-C25 in Guvenen et al. (2021). Block bootstrap standard errors are in parentheses.
C.3 Sources of tail income shocks
Table C.3 provides evidence on the sources of large income reductions that contribute to income left skewness and excess kurtosis. Columns 1 & 2 report average unemployment, disability, bad health, etc, in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively. Column 3 reports the $p$-value of the difference between the two; in most cases, the bottom 10% (effectively households facing large income reductions) experience unemployment, worsening health, large drops in hours, and precarious housing at much higher rates than the rest.